H.W 13 Q5 QUESTION 5 The manager of a travel agency asked you to come up with a forecasting technique that will best fit to the actual
Answer #1

Step 1:

Step 2:

F1 Technique:

CFE = -40

MAD = 251/8 = 31.38

F2 Technique:

CFE = 124

MAD = 280.45/8 = 35.06

F1 is the best